# AUTHOR ONE et al

### v1

@@ Abstract
[Summary]We revisit the question of what randomization ratio (RR) maximizes power of the logrank test in event-driven survival trials under proportional hazards (PH). By comparing three approximations of the logrank test (Schoenfeld, Freedman, Rubinstein) to empirical simulations, we find that the RR that maximizes power is the RR that balances number of events across treatment arms at the end of the trial. This contradicts the common misconception implied by Schoenfeld’s approximation that 1:1 randomization maximizes power. Besides power, we consider other factors that might influence the choice of RR (accrual, trial duration, sample size, etc.). We perform simulations to better understand how unequal randomization might impact these factors in practice. Altogether, we derive 6 insights to guide statisticians in the design of survival trials considering unequal randomization.
@@ 1 Introduction
Practice 2 can be attributed to Schoenfeld1, whose approximation for the large-sample distribution of the logrank test established number of events as the effective sample size under proportional hazards (PH). Practice 1, on the other hand, may involve multiple considerations, including patient preference, cost, ethics, and statistical power.2, 3, 4, 5, 6 With respect to power, it is often believed that 1:1 randomization maximizes power of the logrank test under PH. From the authors’ collective experience, this belief has been perpetuated by academic, industry, and government statisticians alike. This belief is however unfounded. Randomizing patients 1:1 does not in fact maximize power in event-driven trials under PH.
@@
This paper seeks to rectify the common misconception that 1:1 randomization maximizes power of the logrank test in event-driven survival trials under PH. To that end, it is helpful to consider why this misconception exists to begin with. Perhaps it is because 1:1 randomization maximizes power in other settings, e.g., continuous endpoint with common variance. Or because Schoenfeld’s approximation for the logrank test implies that 1:1 randomization maximizes power under PH. We shall later show that Schoenfeld’s approximation increasingly underestimates power when more patients are randomized to the experimental arm than the control arm. Of course, an equation that is biased in favor of 1:1 randomization cannot itself be used to justify that 1:1 randomization maximizes power.
@@
Besides Schoenfeld, other approximations for the logrank test under PH have been proposed which also have immediate, albeit different, implications for randomization. Sposto used Rubinstein’s approximation for the logrank test to suggest that power can be gained by randomizing more patients to the experimental arm.7, 8 Hsieh used Freedman’s approximation to suggest that power is maximized when the randomization ratio is equal to 1/(hazard ratio).9, 10
@@
But despite these previous publications, a question remains unclear. What randomization ratio (RR) maximizes power in event-driven trials if it isn’t 1:1? Sposto considered time-driven trials (i.e., trials in which patients are followed for a fixed period of time), which is different from common practice (Practice 2). Hsieh only compared Freedman’s approximation to Schoenfeld’s. Freedman’s approximation could thus face the same challenge as Schoenfeld’s. That is, if it is found to have differential accuracy under unequal randomization, then its conclusion of when power is maximized is equally questionable.
@@
The rest of the article is organized as follows. In Section 2, we review Schoenfeld’s, Freedman’s, and Rubinstein’s approximation for the logrank test and their implications on randomization in event-driven trials under PH. We then compare their power calculations to empirical simulations over a wide range of scenarios. Our simulation study will reveal that Rubinstein’s approximation is the most accurate of the three, and that it consistently estimates empirical power within 1% absolute error. This will lead us to the conclusion that, for event-driven trials under PH, the RR that maximizes power of the logrank test is the RR that balances number of events across treatment arms at the end of the trial.
@@ 2 Large-sample distribution of the logrank statistic
To assess power of the logrank test, one needs the distribution of the logrank statistic under the alternative hypothesis that treatment is effective. For that purpose, a number of large-sample approximations have been proposed under PH (i.e., hazard ratio is equal to some constant $e^{\theta}=\log S_{e}(t)/\log S_{c}(t)$ for all $t$) which take the form of a normal distribution with some mean $\mu$ and unit variance. Under such normal approximations, the $\phi$ which maximizes statistical power is equivalent to that which minimizes $\mu$.
Next, we review several $\mu$’s and their implied $\operatorname*{argmin}_{\phi}\mu$. Before doing so, it is worth noting that these approximations were originally proposed under the setting that patients are followed until a certain amount of time has elapsed. Thus, some $\mu$’s may in fact depend on trial duration. But as explained by Yung and Liu11, these approximations can be applied to event-driven trials by approximating trial duration with $t_{d}=\inf\{t:E(D(t))=d\}$, the earliest time at which the expected number of accumulated events $E(D(t))$ equals $d$.
@@ 2.1 Schoenfeld (1981)
Under local alternatives (i.e., an alternative that converges to $H_{0}$ as $n$ increases), Schoenfeld1 approximated the large-sample mean of the logrank statistic by
@@
This result is somewhat odd, given that an assumption of local alternatives is used to derive it, but the end result involves a fixed alternative treatment effect $\theta$. It is, however, also elegant. $\mu_{S}$ depends on $S_{e}(\cdot)$ and $S_{c}(\cdot)$ only through the log hazard ratio $\theta$. It does not change regardless of the censoring distribution $L(\cdot)$ or accrual distribution $A(\cdot)$. The fact that $\mu_{S}$ depends on $d$, but not $n$ or $t_{d}$, gives investigators the flexibility to design event-driven trials that are fast (large $n$, small $t_{d}$) or slow (small $n$, large $t_{d}$) depending on the patient population and available resources.12
@@ 2.3 Rubinstein (1981)
Suppose $1-S_{e}\sim\mbox{Exp}(\lambda_{e})$ and $1-S_{c}\sim\mbox{Exp}(\lambda_{c})$. That is, events in each arm follows an exponential distribution. Rubinstein7 derived the maximum likelihood estimates (MLEs) for the event hazard rates and the corresponding test for $H_{0}$: $\log(\lambda_{e}/\lambda_{c})=0$. Their wald test based on the exponential MLEs can be used to approximate the logrank test under the argument that the latter is asymptotically efficient when there is general noninformative censoring.13 Doing so, we obtain
@@
To gain some intuition as to how this conclusion compares to Schoenfeld’s and Freedman’s, suppose $A\sim\mbox{Uni}(r)$ and $L\sim\mbox{Exp}(\eta)$. That is, accrual and dropout follow uniform and exponential distributions. We show in Appendix A that
@@ 2.4 Which approximation is the most accurate?
We now consider which of the three approximations—Schoenfeld, Freedman, or Rubinstein—is the most accurate for estimating power and for comparing power between different randomization ratios. To that end, suppose that survival in the control arm follows an exponential distribution with median of 12 months and that dropout follows an exponential distribution such that patients have a 1% probability of dropping out every 12 months. We consider event-driven trials with varying treatment effect and trial duration by defining a grid of values for the hazard ratio (HR) and event-patient ratio $(d/n)$: $(\text{HR},d/n)\in[0.5,0.8]\times[0.5,0.8]$. For each unique pair $(\text{HR},d/n)$, we set
@@
  • accrual rate $A^{\prime}(\cdot)=20+(50-20)(\text{HR}-0.5)/(0.8-0.5)$, which increases from 20 to 50 patients/month as HR increases so that the expected total trial duration is always less than 5 years.
We then calculate power $\text{Power}(\text{RR},\text{M};\text{HR},d/n)$ for different randomization ratios RR (1:1, 3:2, 2:1) using different methods M (S = Schoenfeld, F = Freedman, R = Rubinstein, E = Empirical simulations).
Figure 1: Bias of power calculations comparing various approximations to empirical simulations across scenarios with different event-patient ratios ($d/n$), hazard ratios (HR) and randomization ratios (RR). LOESS smoothing was applied to more clearly display trends amidst Monte Carlo error. The Monte Carlo standard error is estimated to be $\sim$0.6given 5000 simulations and true power $\sim$80%. Note that remaining minor fluctuations in the curves is to be expected, since the x-axis may not exactly represent the actual $d/n$ for a specific scenario; actual $d/n$ may differ slightly from the desired $d/n$ (x-axis) since it also depends on the RR.
@@
$\text{Bias}=\text{Power}(\text{RR},\text{M};\text{HR},d/n)-\text{Power}(\text{RR},\text{E};\text{HR},d/n).$ (9)
As noted by Barthel et al14, Schoenfeld performs well under 1:1 randomization even for HRs as small as 0.5. However, the authors did not consider performance under unequal randomization. Here, we see that Schoenfeld increasingly underestimates empirical power under 3:2 and 2:1 randomization, particularly for effective treatments and fast trials (small HR and $d/n$). For example, when RR = 2:1 and HR = $d/n$ = 0.5, Schoenfeld underestimates power by more than 4%. Freedman may also lead to over- or under-estimation by up to 3% under certain scenarios. Only Rubinstein consistently estimates empirical power accurately within 1%.
@@
Schoenfeld suggests that randomizing patients 2:1 instead of 1:1 decreases power by 4.8%, regardless of HR or $d/n$. Meanwhile, Freedman suggests that difference in power depends on the HR, with 2:1 randomization becoming increasingly favorable as HR decreases (e.g., when HR=0.5, 2:1 randomization increases power by 4.6%). Rubinstein suggests that difference in power depends on both HR and $d/n$. Compared to Freedman, the potential for increase in power is dampened. Only when both HR and $d/n$ are small (<0.6) does 2:1 randomization lead to a marginal increase in power (0-2%). Similar observations can be made comparing 3:2 to 1:1 randomization (Figure S1).
@@
Figure 2: Difference in power between 2:1 and 1:1 randomization, according to various approximations or empirical simulations and across scenarios with different hazard ratios (HR) and event-patient ratios ($d/n$). For difference in power between 3:2 and 1:1 randomization, see Figure S1.
@@
Having compared Schoenfeld’s, Freedman’s, and Rubinstein’s approximations to empirical simulations across many scenarios, it is clear that Rubinstein’s is the most accurate of the three. It consistently estimates power well (Figure 1) and captures the impact of HR and $d/n$ on power for unequal vs. equal randomization most accurately (Figure 2). One only needs to look at specific scenarios to confirm that the randomization ratio that maximizes power is indeed the randomization ratio that balances number of events across treatment arms at the end of the trial—as suggested by Rubinstein’s approximation. Figure 3 illustrates this in four scenarios. (For more scenarios, see Figure S2.)
@@
Figure 3: Power for various randomization ratios, according to 10,000 empirical simulations and Rubinstein’s approximation. Four scenarios with varying hazard ratios (HR) and event-patient ratio $d/n=0.7$ are considered. Black dots indicate the point at which power is maximized under Rubinstein’s approximation. They also indicate the randomization ratio under which events are expected to be balanced across treatment arms at the end of the trial. LOESS smoothing was applied to the empirical curves to more clearly display trends amidst Monte Carlo error. The Monte Carlo standard error is estimated to be $\sim$0.4given 10,000 simulations and true power $\sim$80%.
@@
This conclusion that “power is maximized when events are balanced across arms” is not completely new. Hsieh10 hypothesized that “the optimal design may exist in the neighborhood from equal allocation of sample sizes to equal allocation of events”. Later, Yung and Liu11 narrowed it down to when “events are balanced across arms”. However, Yung and Liu drew their conclusion from observing the application of another normal approximation for the logrank statistic. This approximation—while more flexible because it can be used under arbitrary continuous accrual, survival, and dropout distributions—cannot be rewritten in such a way as Rubinstein’s to arrive at Equation (6). Therefore, the current work can be seen as providing a theoretical proof to substantiate Yung and Liu’s observation.
Two more points deserve to be made here. First, Figure 2 indicates that randomizing patients 2:1 instead of 1:1 may lead to some loss of power, but not to the level implied by Schoenfeld’s approximation. Investigators should therefore choose carefully which approximation they use when comparing power between different randomization ratios. Whereas Schoenfeld and Freedman may exaggerate the loss or gain in power, Rubinstein and empirical simulations provide more accurate power calculations.
@@
Second, we see from Figure 3 that while 1:1 randomization may not maximize power, its power is often close to the randomizatio ratio that does. Only in the most extreme scenario (HR $=d/n=0.5$) does 1:1 randomization lead to a 2% loss of power. This observation is consistent with Sposto’s conclusion for time-driven trials: “Although one can compute the allocation which maximizes power, the resulting increase in power would be inconsequential. Rather, the fact that a range of allocations yields practically the same power allows one to consider unequal allocations when they would be desirable for other reasons”.8 This ties nicely into Section 3 where we consider some of these “other reasons”.
2.5 Generalizing Rubinstein’s Approximation
@@
Rubinstein’s approximation was derived under the assumption of exponential survival. Its accuracy was also illustrated in the previous subsection under exponential survival. Therefore, one might wonder what happens when the PH assumption is true, but survival does not follow an exponential distribution. Can we still conclude that power of the logrank test is maximized by balancing the expected number of events across arms?
@@
To that end, Wu15 considered Weibull survival distributions under PH and showed that the asymptotic mean of the parametric test statistic is equal to $\mu_{R}$ times the common shape parameter. Thus, our conclusion is at least robust to the Weibull distribution. Next, we consider piecewise exponential survival distributions under PH. Piecewise exponential distributions are flexible in that they can approximate almost any survival distribution well with a sufficient number of knots. Using standard MLE theory, we show in Appendix B.1 that the logrank statistic follows approximately a normal distribution with variance 1 and mean
$\mu_{PE}=\theta\left(\sum_{j=1}^{J}\left(\frac{1}{D_{e,j}}+\frac{1}{D_{c,j}}\right)^{-1}\right)^{1/2},$ (10)
where $J-1$ denotes the number of knots and $D_{e,j}$ and $D_{c,j}$ denote the total number of events observed in the experimental arm and control arm during the $j$th time interval of follow-up, respectively.
The ratios $D_{e,j}/D_{c,j}$, $j=1,...,J$ of observed events per time interval will generally differ. Thus, one can no longer easily argue that $\mu_{PE}$ is minimized (with respect to $\phi$) precisely when the total number of events is balanced across arms. Nevertheless, Equation (10) shows that any time interval with a highly imbalanced ratio of events will contribute little information. One can therefore argue, at least heuristically, that $\mu_{PE}$ is minimized when the number of events contributed by each arm is well-balanced over the multiple time intervals, especially intervals with more events.
We looked at numerous event-driven trials with varying piecewise exponential hazards ($J=2$). In every case, $\mu_{PE}$ was minimized when the expected number of events was approximately equal across treatment arms (see Appendix B.2 for more details). Thus, it would seem that our conclusion “power is maximized when events are balanced across arms” is not restricted to exponential survival distributions, but is rather applicable to general settings under PH.

### v2

@@ Abstract
[Summary]We revisit the question of what randomization ratio (RR) maximizes power of the logrank test (LRT) in event-driven survival trials under proportional hazards (PH). By comparing three approximations of the LRT (Schoenfeld, Freedman, and Rubinstein) to empirical simulations, we find that the RR that maximizes power is the RR that balances number of events across treatment arms at the end of the trial. This contradicts the common misconception implied by Schoenfeld’s approximation that 1:1 randomization maximizes power. Besides power, we consider other factors that might influence the choice of RR (accrual, trial duration, sample size, etc.). We perform simulations to better understand how unequal randomization might impact these factors in practice. Altogether, we derive 5 insights to guide statisticians in the design of survival trials considering unequal randomization.
@@ 1 Introduction
Practice 2 can be attributed to Schoenfeld1, whose approximation for the large-sample distribution of the logrank test (LRT) established number of events as the effective sample size under proportional hazards (PH). Practice 1, on the other hand, may involve multiple considerations, including patient preference, cost, ethics, and statistical power.2, 3, 4, 5, 6 With respect to power, it is often believed that 1:1 randomization maximizes power of the LRT under PH. From the authors’ collective experience, this belief has been perpetuated by academic, industry, and government statisticians alike. This belief is however unfounded. Randomizing patients 1:1 does not in fact maximize power in event-driven trials under PH.
@@
This paper seeks to rectify the common misconception that 1:1 randomization maximizes power of the LRT in event-driven survival trials under PH. To that end, it is helpful to consider why this misconception exists to begin with. Perhaps it is because 1:1 randomization maximizes power in other settings, e.g., continuous endpoint with common variance. Or because Schoenfeld’s approximation for the LRT implies that 1:1 maximizes power under PH. We shall later show that Schoenfeld’s approximation increasingly underestimates power when more patients are randomized to the experimental arm than the control arm. Of course, an equation that is biased in favor of 1:1 cannot itself be used to justify that 1:1 randomization maximizes power.
@@
Besides Schoenfeld, other approximations for the LRT under PH have been proposed which also have immediate, albeit different, implications for randomization. Sposto used Rubinstein’s approximation for the LRT to suggest that power can be gained by randomizing more patients to the experimental arm.7, 8 Hsieh used Freedman’s approximation to suggest that power is maximized when the randomization ratio is equal to 1/(hazard ratio).9, 10
@@
But despite these previous publications, a question remains unclear. What randomization ratio (RR) maximizes power in event-driven trials if it isn’t 1:1? Sposto considered time-driven trials (i.e., trials in which patients are followed for a fixed period of time), which is different from common practice (Practice 2). Hsieh only compared Freedman’s approximation to Schoenfeld’s. Freedman’s approximation could thus face the same challenge as Schoenfeld’s. That is, if it is found to have differential accuracy under unequal randomization, then its conclusion of when power is maximized is equally questionable. Abel, Jensen, et al.11 identified seven issues related to the assumptions, applicability, and practical use of Schoenfeld and Freedman’s approximations. However, randomization ratio was not one of the topics explored by the authors.
@@
The rest of the article is organized as follows. In Section 2, we review Schoenfeld’s, Freedman’s, and Rubinstein’s approximations for the LRT and their implications on randomization in event-driven trials under PH. We then compare their power calculations to empirical simulations over a wide range of scenarios. Our simulation study will reveal that Rubinstein’s approximation is the most accurate of the three, and that it consistently estimates empirical power within 1% absolute error. This will lead us to the conclusion that, for event-driven trials under PH, the RR that maximizes power of the LRT is the RR that balances number of events across treatment arms at the end of the trial. A helpful way to remember this is to remember to balance the effective sample size across treatment arms. For continuous endpoints, that means balancing number of patients. For time-to-event endpoints, that means balancing number of events.
@@ 2 Large-sample distribution of the logrank statistic
To assess power of the LRT, one needs the distribution of the logrank statistic under the alternative hypothesis that treatment is effective. For that purpose, a number of large-sample approximations have been proposed under PH (i.e., hazard ratio is equal to some constant $e^{\theta}=\log S_{e}(t)/\log S_{c}(t)$ for all $t$) which take the form of a normal distribution with some mean $\mu$ and unit variance. Under such normal approximations, the $\phi$ which maximizes statistical power is equivalent to that which minimizes $\mu$.
Next, we review several $\mu$’s and their implied $\operatorname*{argmin}_{\phi}\mu$. Before doing so, it is worth noting that these approximations were originally proposed under the setting that patients are followed until a certain amount of time has elapsed. Thus, some $\mu$’s may in fact depend on trial duration. But as explained by Yung and Liu12, these approximations can be applied to event-driven trials by approximating trial duration with $t_{d}=\inf\{t:E(D(t))=d\}$, the earliest time at which the expected number of accumulated events $E(D(t))$ equals $d$.
@@ 2.1 Schoenfeld (1981)
Under local alternatives (i.e., an alternative that converges to $H_{0}$ as $n$ increases) and assuming that censoring due to loss of follow-up is independent of survival and equal across treatment arms, Schoenfeld1 approximated the large-sample mean of the logrank statistic by
@@
This result is somewhat contradicting, given that two different treatment effects (local alternatives and fixed treatment effect $\theta$) were assumed to approximate/simplify different parts of LRT’s asymptotic distribution. It is, however, also elegant. $\mu_{S}$ depends on $S_{e}(\cdot)$ and $S_{c}(\cdot)$ only through the log hazard ratio $\theta$. It does not change regardless of the censoring or accrual distributions. The fact that $\mu_{S}$ depends on $d$, but not $n$ or $t_{d}$, gives investigators the flexibility to design event-driven trials that are fast (large $n$, small $t_{d}$) or slow (small $n$, large $t_{d}$) depending on the patient population and available resources.13
@@ 2.3 Rubinstein (1981)
Suppose events in the experimental and control arms follow exponential distributions with hazard rates $\lambda_{e}$ and $\lambda_{c}$, respectively. Under arbitrary accrual distribution and arbitrary censoring distributions that are independent of survival but not necessarily equal across treatment arms, Rubinstein7 derived the maximum likelihood estimates (MLEs) for the hazard rates and the corresponding test for $H_{0}$: $\log(\lambda_{e}/\lambda_{c})=0$. Their wald test based on the exponential MLEs can be used to approximate the LRT under the argument that the LRT is asymptotically efficient when there is general noninformative censoring.14 Doing so, we obtain
@@
To gain some intuition as to how this conclusion compares to Schoenfeld’s and Freedman’s, suppose that accrual follows a uniform distribution and that dropout in both arms follows an exponential distribution with rate $\eta$. We show in Appendix A that
@@ 2.4 Which approximation is the most accurate?
We now consider which of the three approximations—Schoenfeld, Freedman, or Rubinstein—is the most accurate for estimating power and for comparing power between different randomization ratios. To that end, suppose that survival in the control arm follows an exponential distribution with median of 12 months and that dropout follows an exponential distribution such that patients have a 1% probability of dropping out every 12 months. We consider event-driven trials with varying treatment effect and trial duration by defining a grid of values for the hazard ratio (HR) and event-patient ratio (EPR $=d/n$): $(\text{HR},\text{EPR})\in[0.5,0.8]\times[0.5,0.8]$. For each unique pair (HR, EPR), we set
@@
  • accrual rate $20+(50-20)(\text{HR}-0.5)/(0.8-0.5)$, which increases from 20 to 50 patients/month as HR increases so that the expected total trial duration is always less than 5 years.
[…]
Figure 1: Bias of power calculations comparing various approximations to empirical simulations across scenarios with different event-patient ratios (EPR), hazard ratios (HR) and randomization ratios (RR). LOESS smoothing was applied to more clearly display trends amidst Monte Carlo error. The Monte Carlo standard error is estimated to be $\sim$0.6given 5000 simulations and true power $\sim$80%. Note that remaining minor fluctuations in the curves is to be expected, since the x-axis may not exactly represent the actual EPR for a specific scenario; actual EPR may differ slightly from the desired EPR (x-axis) since it also depends on the RR.
@@
$\text{Bias}=\text{Power}(\text{RR},\text{M};\text{HR},\text{EPR})-\text{Power}(\text{RR},\text{E};\text{HR},\text{EPR}).$ (9)
As noted by Barthel et al15, Schoenfeld performs well under 1:1 even for HRs as small as 0.5. Although the authors did not consider performance under unequal randomization, here we see that Schoenfeld increasingly underestimates empirical power under 3:2 and 2:1 randomization, particularly for effective treatments and fast trials (small HR and EPR). For example, when RR = 2:1 and HR = EPR = 0.5, Schoenfeld underestimates power by more than 4%. Freedman may also lead to over- or under-estimation by up to 3% under certain scenarios. Only Rubinstein consistently estimates empirical power accurately within 1%.
@@
Schoenfeld suggests that randomizing patients 2:1 instead of 1:1 decreases power by 4.8%, regardless of HR or EPR. Meanwhile, Freedman suggests that difference in power depends on the HR, with 2:1 randomization becoming increasingly favorable as HR decreases (e.g., when HR=0.5, 2:1 randomization increases power by 4.6%). Rubinstein suggests that difference in power depends on both HR and EPR. Compared to Freedman, the potential for increase in power is dampened. Only when both HR and EPR are small (<0.6) does 2:1 randomization lead to a marginal increase in power (0-2%). Similar observations can be made comparing 3:2 to 1:1 randomization (Figure S1).
@@
Figure 2: Difference in power between 2:1 and 1:1 randomization, according to various approximations or empirical simulations and across scenarios with different hazard ratios (HR) and event-patient ratios (EPR). For difference in power between 3:2 and 1:1 randomization, see Figure S1.
@@
Having compared Schoenfeld’s, Freedman’s, and Rubinstein’s approximations to empirical simulations across many scenarios, it is clear that Rubinstein’s is the most accurate of the three. It consistently estimates power well (Figure 1) and captures the impact of HR and EPR on power for unequal vs. equal randomization most accurately (Figure 2). One only needs to look at specific scenarios to confirm that the randomization ratio that maximizes power is indeed the randomization ratio that balances number of events across treatment arms at the end of the trial—as suggested by Rubinstein’s approximation. Figure 3 illustrates this in four scenarios. (For more scenarios, see Figure S2.)
@@
Figure 3: Power for various randomization ratios, according to 10,000 empirical simulations and Rubinstein’s approximation. Four scenarios with varying hazard ratios (HR) and event-patient ratio $0.7$ are considered. Black dots indicate the point at which power is maximized under Rubinstein’s approximation. They also indicate the randomization ratio under which events are expected to be balanced across treatment arms at the end of the trial. LOESS smoothing was applied to the empirical curves to more clearly display trends amidst Monte Carlo error. The Monte Carlo standard error is estimated to be $\sim$0.4given 10,000 simulations and true power $\sim$80%.
@@
This conclusion that “power is maximized when events are balanced across arms” is not completely new. Hsieh10 hypothesized that “the optimal design may exist in the neighborhood from equal allocation of sample sizes to equal allocation of events”. Later, Yung and Liu12 narrowed it down to when “events are balanced across arms”. However, Yung and Liu drew their conclusion from observing the application of another normal approximation for the logrank statistic. This approximation—while more flexible because it can be used under arbitrary survival distributions—cannot be rewritten in such a way as Rubinstein’s to arrive at Equation (6). Therefore, the current work can be seen as providing further substantiation of Yung and Liu’s observation.
@@
While 1:1 may not maximize power, we see from Figure 3 that its power is often close to the randomization ratio that does. Only in the most extreme scenario (HR = EPR = 0.5) does 1:1 randomization lead to a 2% loss of power. This observation is consistent with Sposto’s conclusion for time-driven trials: “Although one can compute the allocation which maximizes power, the resulting increase in power would be inconsequential. Rather, the fact that a range of allocations yields practically the same power allows one to consider unequal allocations when they would be desirable for other reasons”.8 This ties nicely into Section 3 where we consider some of these “other reasons”.
2.5 Generalizing Rubinstein’s Conclusion
@@
Rubinstein’s approximation was derived under exponential survival, arbitrary accrual, and arbitrary censoring distributions. Its accuracy was further illustrated in the previous subsection under exponential survival. Therefore, one might wonder what happens under general PH. Can we still conclude that power of the LRT is maximized by balancing the expected number of events across arms?
@@
Wu16 considered Weibull survival distributions under PH and showed that the asymptotic mean of the parametric test statistic is equal to $\mu_{R}$ times the common shape parameter. Thus, our conclusion is at least robust to the Weibull distribution.
For the case under general PH, an argument proceeds as follows. Given two survival functions under PH, there exists a monotonic transformation of the survival times such that the transformed observations follow two exponential distributions.17 Rubinstein’s approximation implies that power of the LRT—when applied to the transformed observations—is maximized when events are balanced across arms. Meanwhile, the LRT is a rank-based test, so a monotonic transformation does not change the log-rank statistic itself. The transformation also does not change the total number of events (assuming an event-driven trial) nor the ordering of events across arms. Therefore, if balancing events across arms maximizes power for the transformed observations (i.e., exponential survival curves), then it too maximizes power for the original observations (i.e, survival curves under general PH).